OpenStudy (anonymous):

Find the product. (-4·3·2)^2

OpenStudy (anonymous):

\[(a\cdot b)^c = a^c \cdot b^c\]

OpenStudy (anonymous):

576

OpenStudy (anonymous):

Or you can evaluate the product inside, then square that. Either way.

OpenStudy (anonymous):

(-4·3·2)^2 = (-24)^2 = 576 lol

OpenStudy (anonymous):

Polack....Parentheses first seems to be the simplest approach

OpenStudy (anonymous):

Heromeals... Yeah, but often they are given variables rather than easy to compute numbers, so the general rule is also important.